# What Impulse Response Do Instrumental Variables Identify?

### v3

@@ 3.4.2 Sharpness under Weak Stationarity
$ID(\text{AC}_{P}\;;\>\bm{\Xi})=\Bigl{\{}\Pi(\bm{\tau}):\bm{\tau}\in\bm{\Xi}\>\text{ generates }\text{AC}_{P}\Bigr{\}}.$ (21)
@@
where $\bm{\Gamma}=[\gamma_{1}^{A}\quad\gamma_{2}^{A}\>;\>\gamma_{1}^{B}\quad\gamma_{2}^{B}]$. If we can assume $\bm{\Gamma}=0$, then (27) is empirically testable. Indeed, we note that Jarociński and Karadi (2020) put the same structures on IVs. In their empirical model, the sample correlation coefficient between the IVs is negative (.4261-.4261- .4261,  s.e.=.1391=.1391= .1391), which indicates that the premise of Proposition 4.3 holds, in which case and the marginal identified set for $\theta_{1,y2}/\theta_{0,x2}$ [^11] This corresponds to the impulse response to the central bank information shock in Jarociński and Karadi (2020). is unbounded. This entails that any bounded confidence set has zero coverage probability (Dufour, 1997). Since their procedure always produces a bounded credible set, its frequentist size must also be zero. On the other hand, our LP-IV inference framework accommodates unbounded identified sets, providing valid one-sided confidence intervals as discussed in Appendix C.2.
@@
$\begin{array}[]{ccccccccc}\widehat{\beta}^{RZ}_{h}&=&\widehat{w}_{h,1}^{RZ}&\times&\widehat{\frac{\widetilde{\theta}_{h,y1}}{\widetilde{\psi}_{h,11}}}&+&\widehat{w}_{h,2}^{RZ}&\times&\widehat{\frac{\widetilde{\theta}_{h,y2}}{\widetilde{\psi}_{h,22}}},\\ 0.37&&1.87&&0.68&&-0.87&&1.02\\ \end{array}$ (29)
@@ 5  Monetary Policy and Central Bank Information Shock
Applying the sign restrictions approach from Section 3.3, we use LP-IV estimates as identification bounds for impulse responses to component shocks within monetary policy announcements. Specifically, $z_{t}^{ff}$ gives $\Theta_{++}$ and $z_{t}^{sp}$ gives $\Theta_{+-}$ [^18] We use $-z_{t}^{sp}$ to apply Corollary 2 since $\operatorname*{cov}(z_{t}^{sp},x_{t})<0$. in Corollary 2, with their intersection corresponding to cases (ii) and (iii) in Proposition 3. Thus, for each $h$, the identified set of responses to the monetary policy shock is $\{\theta_{h,y1}:\min(\beta_{h}^{ff},\beta_{h}^{sp})<\theta_{h,y1}<\max(\beta_{h}^{ff},\beta_{h}^{sp})\}$, while for the central bank information shock it is $\{\theta_{h,y2}:\theta_{h,y2}>\beta_{h}^{ff}\}$ if $\beta_{h}^{ff}>\beta_{h}^{sp}$ and $\{\theta_{h,y2}:\theta_{h,y2}<\beta_{h}^{ff}\}$ if $\beta_{h}^{ff}<\beta_{h}^{sp}$.
@@ A.1 Proof of Proposition 1
Since $\theta_{0,xs}>0$ for $s=1,2,...,S$, the statement of the proposition follows. ∎
@@ A.3 Proof of Proposition 2
[…]
@@ A.6 Lemmas for Proposition 4
The following lemma will be used repeatedly in the proof of Proposition 4. For these two lemmas and the proof of Proposition 4, with abuse of notation, we augment the $\times 2$ 2 matrix $\mathbf{D}$ by appending a 2×12$\times$ 12 × 1 column zero vector and still denote this 2×32$\times$ 32 × 3 augmented matrix by $\mathbf{D}$.
@@
Proof: Notice that the autocovariance of $(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}$ of order 222 or greater is a zero matrix:
@@
Proof: 1. Given $\bm{\tau}\in\bm{\Xi}$, define $\widetilde{\bm{\tau}}=(\widetilde{\bm{\Theta}}_{0},\widetilde{\bm{\Theta}}_{1},\widetilde{\bm{\mathbf{\lambda}}},\widetilde{\mathbf{D}},\widetilde{\bm{\Gamma}})^{\prime}$ where $\widetilde{\bm{\Theta}}_{s}=\bm{\Theta}_{s}\text{diag}(\theta_{0,x1}^{-1},\theta_{0,x2}^{-1},1)$ for $s=0,1$, $\widetilde{\bm{\mathbf{\lambda}}}=((\theta_{0,x1})^{2}\lambda_{1},(\theta_{0,x2})^{2}\lambda_{2},\lambda_{3})^{\prime}$, $\widetilde{\mathbf{D}}=\mathbf{D}\text{diag}(\theta_{0,x1}^{-1},\theta_{0,x2}^{-1},1)$, and $\widetilde{\bm{\Gamma}}=\bm{\Gamma}$. Then, $\widetilde{\bm{\tau}}\in\bm{\Xi}$ and $\widetilde{\theta}_{0,x1}=\widetilde{\theta}_{0,x1}=1$ so that $\widetilde{\bm{\tau}}\in\bm{\Xi}^{(\text{EN})}$. Furthermore, for $s,\tilde{s}\in\{0,1\}$,
@@
$\Pi(\widetilde{\bm{\tau}})=$ $\left(\widetilde{\theta}_{1,y1}/\widetilde{\theta}_{0,x1},\widetilde{\theta}_{1,y2}/\widetilde{\theta}_{0,x2}\right)^{\prime}$
@@
2. Given $\bm{\tau}\in\bm{\Xi}$, define $\grave{\bm{\tau}}=(\grave{\bm{\Theta}}_{0},\grave{\bm{\Theta}}_{1},\grave{\bm{\mathbf{\lambda}}},\grave{\mathbf{D}},\grave{\bm{\Gamma}})^{\prime}$ where $\grave{\bm{\Theta}}_{s}=\bm{\Theta}_{s}\text{diag}(\lambda_{1}^{1/2},\lambda_{2}^{1/2},\lambda_{3}^{1/2})$ for $s=0,1$, $\grave{\bm{\mathbf{\lambda}}}=(1,1,1)^{\prime}$, $\grave{\mathbf{D}}=\mathbf{D}\text{diag}(\lambda_{1}^{1/2},\lambda_{2}^{1/2},\lambda_{3}^{1/2})$, and $\widetilde{\bm{\Gamma}}=\bm{\Gamma}$. Then, $\grave{\bm{\tau}}\in\bm{\Xi}^{(\text{VN})}$ and for $s,\tilde{s}=0,1$,
@@
$\Pi(\grave{\bm{\tau}})=$ $\left(\dfrac{\lambda_{1}^{1/2}\theta_{1,y1}}{\lambda_{1}^{1/2}\theta_{0,x1}},\dfrac{\lambda_{2}^{1/2}\theta_{1,y2}}{\lambda_{2}^{1/2}\theta_{0,x2}}\right)^{\prime}$
@@ A.7 Proof of Proposition 4
It suffices to show that the claims hold for $\overline{\bm{\Xi}}=\bm{\Xi}^{(\text{VN})}$ by Lemma A.1. Let $\bm{\tau}\in\bm{\Xi}^{(\text{VN})}$ and $\mathbf{R}_{r}$ be an 222-dimensional rotation/reflection matrix with angle $r$:
@@
[…]
@@
i.e., when $R_{r}$ is a rotation, and the range of $r$ which satisfies these conditions is given by $[0,r^{*})\cup(r^{**},2\pi)$ where $r^{*}$ solves over $r\in[0,\pi/2)$:
@@
Furthermore, observe that both $\overline{\theta}{}_{1,y1}(r)/\overline{\theta}_{0,x1}(r)$ and $\overline{\theta}{}_{1,y2}(r)/\overline{\theta}_{0,x2}(r)$, as functions of $r$, are decreasing in $[0,r^{*})$ and increasing in $(r^{**},2\pi)$.
@@
Then, $(\underline{r}^{(s)},\overline{r}^{(s)})$ are unique over $[0,\pi/2)\cup(3\pi/2,2\pi)$, $\underline{r}^{(s)}\in[0,2\pi),\>\overline{r}^{(s)}\in(3\pi/2,2\pi)$ for $s=1,2$ and, satisfy
@@ B.1 Augmented SVMA
The augmented SVMA model extends the baseline SVMA model (5) to include the disaggregated variables, $x_{s,t}$ for $s=1,2,...,S$ such that $x_{t}=\sum_{s=1}^{S}x_{s,t}$. The $x_{s,t}$’s are the components of the aggregate variable $x_{t}$. The augmented SVMA model is given by
@@
where $\bm{Y}_{t}^{A}=(x_{1,t},x_{2,t},...,x_{S,t},...,y_{t})^{\prime}$ is an $(n+S-1)\times 1$ vector of observed endogenous variables, $\bm{\Psi}(L)=\bm{\Psi}_{0}+\bm{\Psi}_{1}L+\bm{\Psi}_{2}L^{2}+\cdots$, and $\bm{\Psi}_{h}$ for $h=0,1,2,...,$ is an $(n+S-1)\times m$ matrix of impulse responses. Let $\bm{B}=(\bm{i}_{n}~\bm{i}_{n}~\cdots~\bm{i}_{n}~\bm{I}_{n})$ be the $n\times(n+S-1)$ matrix where $\bm{i}_{n}=(1,0,...,0)^{\prime}$ and $\bm{I}_{n}$ is the $n\times n$ identity matrix. By pre-multiplying $\bm{B}$ both sides of (45), we can obtain (5) and $\bm{\Theta}(L)=\bm{B}\bm{\Psi}(L)$. Without any further restrictions on the impulse response matrices, (45) is more general than (5). Note that the last $n-1$ rows of $\bm{\Psi}(L)$ are identical to the last $n-1$ rows of $\bm{\Theta}(L)$.
[…]
@@ C.1 LP-IV Estimator and Standard Error
for $t=1,2,...,T$ and $h=0,1,2,...,H$, where the control variables $\bm{R}_{t}$ include a constant and the lagged values of the endogenous variable and the instrument.
@@ C.2 Confidence Intervals
[…]
@@
where $\bm{a}$ is an $S\times 1$ non-random vector, $\{\xi_{t}=1\}$ is a collection of realizations of $\bm{e}_{t}$ such that $\xi_{t}=1$, and $\bm{\theta}_{h,y}=(\theta_{h,y1},\theta_{h,ys},...,\theta_{h,yS})^{\prime}$. Thus, $\theta_{h,y\xi}$ is a weighted average of component-wise impulse responses and the weights are the mean of $\varepsilon_{s,t}$ conditional on $\xi_{t}=1$, which are non-negative for a large class of reasonable distributions of $\bm{e}_{t}$. For example, if we assume $\varepsilon_{s,t}$ for $s=1,...,S$ are i.i.d., then $E[\varepsilon_{s,t}|\xi_{t}=1]=1/S$ and (A.2) becomes the equal-weighted average. Unlike the ATE, heterogeneity arises in the composition of $\xi_{t}$, so that the treatment is heterogeneous. Since the component-wise impulse responses are assumed constant, the treatment effect of a particular composition in $\xi_{t}$ is homogeneous.
@@
$\text{cov}(z_{t}^{A},x_{t})$ $\text{cov}(z_{t}^{B},x_{t})$ $\alpha_{1}^{A}$ $\alpha_{2}^{A}$ $\alpha_{1}^{B}$ $\alpha_{2}^{B}$ $\beta_{h}^{A}-\beta_{h}^{B}$ $\theta_{h_{y}1}$ $\theta_{h,y2}$ + + + + + + $\cdot$ $UL$ $UL$ + + + -- + -- $\cdot$ + + -- + -- + $\cdot$ + -- + + -- -- $\cdot$ + -- + -- -- + $\cdot$ + -- -- + + -- $\cdot$ -- + + -- -- + $\cdot$ -- + -- + + -- $\cdot$ -- + -- -- + + $\cdot$ -- -- + -- + -- $\cdot$ -- -- -- + -- + $\cdot$ -- -- -- -- -- -- $\cdot$ + + + + + -- + $T$ $L$ + + + -- + + -- + -- + + -- + + + -- + -- -- -- -- -- + -- + + + -- -- + -- -- + -- + -- -- -- + -- -- -- -- -- -- -- -- + + + + + + + -- -- $T$ $U$ + + + -- + + + + -- + + -- + -- + -- + -- -- -- + -- + -- + + + + -- + -- -- + -- -- -- -- -- + -- -- + -- -- -- -- -- + -- + + + + -- + + $L$ $T$ + + -- + + + -- + -- + + + -- + + -- -- + -- -- -- -- + + -- + + -- -- + -- -- -- + + -- -- + -- -- -- -- -- -- -- -- + -- + + + + + -- + -- $U$ $T$ + + -- + + + + + -- + + + -- -- + -- -- + -- -- + -- + + -- + + + -- + -- -- -- + -- -- -- + -- -- -- + -- -- -- -- + -- -- + + + -- -- + $\cdot$ $T$ $T$ + + -- + + -- $\cdot$ + -- + -- + -- $\cdot$ + -- -- + -- + $\cdot$ -- + + -- + -- $\cdot$ -- + -- + -- + $\cdot$ -- -- + -- -- + $\cdot$ -- -- -- + + -- $\cdot$
@@
    $\widetilde{\theta}_{h,xs}>0$ for $h=1,2,...,H$ and $s=1,2$: Both of defense and non-defense spending shocks have positive effects on the cumulative government spending over the $h$ periods.
@@
Figure CA.3: Identified Set for Sectoral Spending Multipliers using Sectoral Spending Data and the RZ News Shock Instrument. $\bullet$: Calibrated defense spending multiplier is transitory; $+$: Calibrated defense spending multiplier is persistent.
Figure CA.3 illustrates the identified set for each $h=0,...,20$ (represented by blue solid lines). As $h$ increases, the lines exhibit steeper slopes with larger y-intercepts. This observation implies that as time progresses, for a given magnitude of the defense spending multiplier (x-axis), the corresponding magnitude of the non-defense spending multiplier becomes larger.
@@
[…]
@@
Analogous to the SVMA(1)(1)( 1 ) model in Section 3.4.1, we define observational equivalence with respect to the autocovariances of $\{(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}\}_{t}$ for the finite-dimensional parameter component $\bm{\Xi}$ defined on the parameter space $\bm{\Xi}$ given by
@@
Within this proof, with abuse of notation, we augment the $\times 2$ 2 matrix $\mathbf{D}$ by appending a $2\times(m-2)$ column zero vector and still denote this $m\times 2$ augmented matrix by $\mathbf{D}$.
@@
We are going to show that if there exists $\widetilde{\bm{\tau}}\in\bm{\Xi}^{(\text{VN})}$ such that $\widetilde{\Gamma}$ is a nonzero matrix and $\widetilde{\bm{\tau}}$ generates the same variance of $\mathbf{Z}_{t}$ and the covariance between $x_{t}$ and $\mathbf{Z}_{t}$ as those of $\bm{\tau}$, then it must be that $(\widetilde{\theta}_{0,x1})^{2}+(\widetilde{\theta}_{0,x2})^{2}>2$ Under $\widetilde{\bm{\tau}}\in\bm{\Xi}$, the variance of $x_{t}$ must be no smaller than $(\widetilde{\theta}_{0,x1})^{2}+(\widetilde{\theta}_{0,x2})^{2}>2$ while it is 222 under $\bm{\tau}\in\bm{\Xi}$, failing to match $\text{var}(x_{t})$.
@@
where $Q$ is a $\times 2$ 2 orthogonal matrix and $(q_{1},q_{2})\in(0,1)^{2}\backslash\{(1,1)\}$. Given (A.20), we have
@@
Now, observe that any $\times 2$ 2 orthogonal matrix can be expressed as either a rotation or reflection matrix:
@@
for some $\times 2$ 2 orthogonal matrix $\mathbf{R}$ and
[…]

### v4

@@ 3.4.2 Sharpness under Weak Stationarity
[…]
@@
where $\bm{\Gamma}=[\gamma_{1}^{A}\quad\gamma_{2}^{A}\>;\>\gamma_{1}^{B}\quad\gamma_{2}^{B}]$. If we can assume $\bm{\Gamma}=0$, then (27) is empirically testable. Indeed, we note that Jarociński and Karadi (2020) put the same structures on IVs. In their empirical model, the sample correlation coefficient between the IVs is negative (.4261-.4261,  s.e.=.1391=.1391), which indicates that the premise of Proposition 4.3 holds, in which case and the marginal identified set for $\theta_{1,y2}/\theta_{0,x2}$ [^10] This corresponds to the impulse response to the central bank information shock in Jarociński and Karadi (2020). is unbounded. This entails that any bounded confidence set has zero coverage probability (Dufour, 1997). Since their procedure always produces a bounded credible set, its frequentist size must also be zero. On the other hand, our LP-IV inference framework accommodates unbounded identified sets, providing valid one-sided confidence intervals as discussed in Appendix C.2.
@@
[…]
@@ 5  Monetary Policy and Central Bank Information Shock
Applying the sign restrictions approach from Section 3.3, we use LP-IV estimates as identification bounds for impulse responses to component shocks within monetary policy announcements. Specifically, $z_{t}^{ff}$ gives $\Theta_{++}$ and $z_{t}^{sp}$ gives $\Theta_{+-}$ [^17] We use $z_{t}^{sp}$ to apply Corollary 2 since $\operatorname*{cov}(z_{t}^{sp},x_{t})<0$. in Corollary 2, with their intersection corresponding to cases (ii) and (iii) in Proposition 3. Thus, for each $h$, the identified set of responses to the monetary policy shock is $\{\theta_{h,y1}:\min(\beta_{h}^{ff},\beta_{h}^{sp})<\theta_{h,y1}<\max(\beta_{h}^{ff},\beta_{h}^{sp})\}$, while for the central bank information shock it is $\{\theta_{h,y2}:\theta_{h,y2}>\beta_{h}^{ff}\}$ if $\beta_{h}^{ff}>\beta_{h}^{sp}$ and $\{\theta_{h,y2}:\theta_{h,y2}<\beta_{h}^{ff}\}$ if $\beta_{h}^{ff}<\beta_{h}^{sp}$.
@@ A.1 Proof of Proposition 1
Since $\theta_{0,xs}>0$ for s=1,2,,Ss=1,2,...,S, the statement of the proposition follows. ∎
@@ A.3 Proof of Proposition 2
where $\psi_{0,sr}=E[x_{s,t}|\varepsilon_{r,t}=1]-E[x_{s,t}|\varepsilon_{r,t}=0]$ and $\{\cdots\}$ is a linear combination of the elements in the curly brackets. By Assumption 1, for $j=A,B$ and $s=1,2$,
@@ A.6 Lemmas for Proposition 4
The following lemma will be used repeatedly in the proof of Proposition 4. For these two lemmas and the proof of Proposition 4, with abuse of notation, we augment the $2\times 2$ matrix $\mathbf{D}$ by appending a $2\times 1$ column zero vector and still denote this $2\times 3$ augmented matrix by $\mathbf{D}$.
@@
Proof: Notice that the autocovariance of $(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}$ of order $2$ or greater is a zero matrix:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ A.7 Proof of Proposition 4
It suffices to show that the claims hold for $\overline{\bm{\Xi}}=\bm{\Xi}^{(\text{VN})}$ by Lemma A.1. Let $\bm{\tau}\in\bm{\Xi}^{(\text{VN})}$ and $\mathbf{R}_{r}$ be an $2$-dimensional rotation/reflection matrix with angle $r$:
@@
and $\widetilde{\mathbf{R}}_{r}=\begin{bmatrix}\mathbf{R}_{r}&\bm{0}_{1\times 2}\>;&\>\bm{0}_{2\times 1}&1\end{bmatrix}$. Then, $\widetilde{\mathbf{R}}_{r}$ is an orthogonal matrix: $\widetilde{\mathbf{R}}_{r}\widetilde{\mathbf{R}}_{r}^{\prime}=I_{3}$ and thus $\overline{\bm{\tau}}_{r}=(\bm{\Theta}_{0}\widetilde{\mathbf{R}}_{r},\bm{\Theta}_{1}\widetilde{\mathbf{R}}_{r},\bm{1}_{3\times 1},\mathbf{D}\widetilde{\mathbf{R}}_{r},\bm{\Gamma})^{\prime}$ produces the same autocovariance of $(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}$ as $\bm{\tau}$ by Lemma A.1.
@@
i.e., when $R_{r}$ is a rotation, and the range of $r$ which satisfies these conditions is given by $0,r^{*})\cup(r^{**},2\pi)$ where $r^{*}$ solves over $r\in[0,\pi/2)$:
@@
Furthermore, observe that both $\overline{\theta}{}_{1,y1}(r)/\overline{\theta}_{0,x1}(r)$ and $\overline{\theta}{}_{1,y2}(r)/\overline{\theta}_{0,x2}(r)$, as functions of $r$, are decreasing in $0,r^{*})$ and increasing in $(r^{**},2\pi)$.
@@
Then, $(\underline{r}^{(s)},\overline{r}^{(s)})$ are unique over $0,\pi/2)\cup(3\pi/2,2\pi)$, $\underline{r}^{(s)}\in[0,2\pi),\>\overline{r}^{(s)}\in(3\pi/2,2\pi)$ for $s=1,2$ and, satisfy
@@ B.1 Augmented SVMA
The augmented SVMA model extends the baseline SVMA model (5) to include the disaggregated variables, $x_{s,t}$ for s=1,2,,Ss=1,2,...,S such that $x_{t}=\sum_{s=1}^{S}x_{s,t}$. The $x_{s,t}$’s are the components of the aggregate variable $x_{t}$. The augmented SVMA model is given by
@@
where $\bm{Y}_{t}^{A}=(x_{1,t},x_{2,t},...,x_{S,t},...,y_{t})^{\prime}$ is an $(n+S-1)\times 1$ vector of observed endogenous variables, $\bm{\Psi}(L)=\bm{\Psi}_{0}+\bm{\Psi}_{1}L+\bm{\Psi}_{2}L^{2}+\cdots$, and $\bm{\Psi}_{h}$ for h=0,1,2,,h=0,1,2,..., is an $(n+S-1)\times m$ matrix of impulse responses. Let $\bm{B}=(\bm{i}_{n}~\bm{i}_{n}~\cdots~\bm{i}_{n}~\bm{I}_{n})$ be the $n\times(n+S-1)$ matrix where $\bm{i}_{n}=(1,0,...,0)^{\prime}$ and $\bm{I}_{n}$ is the $n\times n$ identity matrix. By pre-multiplying $\bm{B}$ both sides of (45), we can obtain (5) and $\bm{\Theta}(L)=\bm{B}\bm{\Psi}(L)$. Without any further restrictions on the impulse response matrices, (45) is more general than (5). Note that the last n1n-1 rows of $\bm{\Psi}(L)$ are identical to the last n1n-1 rows of $\bm{\Theta}(L)$.
Let $\psi_{h,rs}$ be the response of disaggregated variable $x_{r,t+h}$ to the component-wise shock $\varepsilon_{s,t}$. This impulse response is the $(r,s)$-th element of $\bm{\Psi}_{h}$. With the unit effect normalization for each component-wise shock ($\psi_{0,ss}=1$ for s=1,2,,Ss=1,2,...,S), $\varepsilon_{s,t}$ is scaled so that one unit change in $\varepsilon_{s,t}$ corresponds to the unit change in $x_{s,t}$. Since $x_{t}=\sum_{s=1}^{S}x_{s,t}$, we write $\sum_{r=1}^{S}\psi_{h,rs}=\theta_{h,xs}$ for s=1,,Ss=1,...,S. That is, the impulse response of $\varepsilon_{s,t}$ on $x_{t}$ is the sum of the impulse responses of $\varepsilon_{s,t}$ on $x_{s,t}$.
@@ C.1 LP-IV Estimator and Standard Error
for t=1,2,,Tt=1,2,...,T and h=0,1,2,,Hh=0,1,2,...,H, where the control variables $\bm{R}_{t}$ include a constant and the lagged values of the endogenous variable and the instrument.
@@ C.2 Confidence Intervals
Let $\widehat{\bm{\beta}}$ be the collection of two IV estimates and suppose that $\sqrt{T}\left(\widehat{\bm{\beta}}-\bm{\beta}\right)\xrightarrow{d}N\left(0,\Sigma\right)$ and $\widehat{\Sigma}$ is a consistent estimator of the asymptotic variance $\Sigma$. The sample size is $T$. For a bivariate random vector $X\sim N\left(0,\Sigma\right)$, let
@@
where $\bm{a}$ is an $S\times 1$ non-random vector, $\{\xi_{t}=1\}$ is a collection of realizations of $\bm{e}_{t}$ such that $\xi_{t}=1$, and $\bm{\theta}_{h,y}=(\theta_{h,y1},\theta_{h,ys},...,\theta_{h,yS})^{\prime}$. Thus, $\theta_{h,y\xi}$ is a weighted average of component-wise impulse responses and the weights are the mean of $\varepsilon_{s,t}$ conditional on $\xi_{t}=1$, which are non-negative for a large class of reasonable distributions of $\bm{e}_{t}$. For example, if we assume $\varepsilon_{s,t}$ for s=1,,Ss=1,...,S are i.i.d., then $E[\varepsilon_{s,t}|\xi_{t}=1]=1/S$ and (A.2) becomes the equal-weighted average. Unlike the ATE, heterogeneity arises in the composition of $\xi_{t}$, so that the treatment is heterogeneous. Since the component-wise impulse responses are assumed constant, the treatment effect of a particular composition in $\xi_{t}$ is homogeneous.
@@
$\text{cov}(z_{t}^{A},x_{t})$ $\text{cov}(z_{t}^{B},x_{t})$ $\alpha_{1}^{A}$ $\alpha_{2}^{A}$ $\alpha_{1}^{B}$ $\alpha_{2}^{B}$ $\beta_{h}^{A}-\beta_{h}^{B}$ $\theta_{h_{y}1}$ $\theta_{h,y2}$ + + + + + + $\cdot$ ULUL ULUL + + + - + - $\cdot$ + + - + - + $\cdot$ + - + + - - $\cdot$ + - + - - + $\cdot$ + - - + + - $\cdot$ - + + - - + $\cdot$ - + - + + - $\cdot$ - + - - + + $\cdot$ - - + - + - $\cdot$ - - - + - + $\cdot$ - - - - - - $\cdot$ + + + + + - + $T$ $L$ + + + - + + - + - + + - + + + - + - - - - - + - + + + - - + - - + - + - - - + - - - - - - - - + + + + + + + - - $T$ $U$ + + + - + + + + - + + - + - + - + - - - + - + - + + + + - + - - + - - - - - + - - + - - - - - + - + + + + - + + $L$ $T$ + + - + + + - + - + + + - + + - - + - - - - + + - + + - - + - - - + + - - + - - - - - - - - + - + + + + + - + - $U$ $T$ + + - + + + + + - + + + - - + - - + - - + - + + - + + + - + - - - + - - - + - - - + - - - - + - - + + + - - + $\cdot$ $T$ $T$ + + - + + - $\cdot$ + - + - + - $\cdot$ + - - + - + $\cdot$ - + + - + - $\cdot$ - + - + - + $\cdot$ - - + - - + $\cdot$ - - - + + - $\cdot$
@@
    $\widetilde{\theta}_{h,xs}>0$ for h=1,2,,Hh=1,2,...,H and $s=1,2$: Both of defense and non-defense spending shocks have positive effects on the cumulative government spending over the $h$ periods.
@@
Figure CA.3: Identified Set for Sectoral Spending Multipliers using Sectoral Spending Data and the RZ News Shock Instrument. $\bullet$: Calibrated defense spending multiplier is transitory; ++: Calibrated defense spending multiplier is persistent.
Figure CA.3 illustrates the identified set for each h=0,,20h=0,...,20 (represented by blue solid lines). As $h$ increases, the lines exhibit steeper slopes with larger y-intercepts. This observation implies that as time progresses, for a given magnitude of the defense spending multiplier (x-axis), the corresponding magnitude of the non-defense spending multiplier becomes larger.
@@
Let $q\geq 1,\>n\geq 2,m\geq 3,L\geq 0$. We consider a general SVMA(q+1)(q+1) model for the pair of $n\times 1$ endogenous and $2\times 1$ instrument vectors $\mathbf{Y}_{t}$ and $\mathbf{Z}_{t}$ with respect to the innovation vector $(\bm{\varepsilon}_{t}^{\prime},\mathbf{v}_{t}^{\prime})^{\prime}$ where $\bm{\varepsilon}_{t}$ is a $m\times 1$ structural shock vector and $\mathbf{v}_{t}$ is a $2\times 1$ instrument specific error vector:
@@
Analogous to the SVMA(1)(1) model in Section 3.4.1, we define observational equivalence with respect to the autocovariances of $\{(\mathbf{Y}_{t}^{\prime},\mathbf{Z}_{t}^{\prime})^{\prime}\}_{t}$ for the finite-dimensional parameter component $\bm{\Xi}$ defined on the parameter space $\bm{\Xi}$ given by
@@
Within this proof, with abuse of notation, we augment the $2\times 2$ matrix $\mathbf{D}$ by appending a $2\times(m-2)$ column zero vector and still denote this $m\times 2$ augmented matrix by $\mathbf{D}$.
@@
We are going to show that if there exists $\widetilde{\bm{\tau}}\in\bm{\Xi}^{(\text{VN})}$ such that $\widetilde{\Gamma}$ is a nonzero matrix and $\widetilde{\bm{\tau}}$ generates the same variance of $\mathbf{Z}_{t}$ and the covariance between $x_{t}$ and $\mathbf{Z}_{t}$ as those of $\bm{\tau}$, then it must be that $(\widetilde{\theta}_{0,x1})^{2}+(\widetilde{\theta}_{0,x2})^{2}>2$ . Under $\widetilde{\bm{\tau}}\in\bm{\Xi}$, the variance of $x_{t}$ must be no smaller than $(\widetilde{\theta}_{0,x1})^{2}+(\widetilde{\theta}_{0,x2})^{2}>2$ while it is $2$ under $\bm{\tau}\in\bm{\Xi}$, failing to match $\text{var}(x_{t})$.
@@
[…]
@@
Now, observe that any $2\times 2$ orthogonal matrix can be expressed as either a rotation or reflection matrix:
@@
for some $2\times 2$ orthogonal matrix $\mathbf{R}$ and
$\widetilde{\mathbf{R}}_{r}=\begin{bmatrix}\mathbf{R}_{r}&\bm{0}_{1\times 2}\>;&\>\bm{0}_{2\times 1}&1\end{bmatrix}$.